# Kalman Filter Enhanced GRPO for Reinforcement Learning-Based Language Model Reasoning

### v1

@@ 5.4.3 KL Divergence, Normalized Gradients and Reward Baseline Stabilization Analysis
(c) The Kalman filtered gradually stabilize the returns to a certain level.
@@
We also examine the KL divergence of action probabilities between RL finetuned model with its reference model as Fig. 4(a). Interestingly, we notice that the KRPO has slightly higher KL divergence compared to GRPO baseline. It makes sence as we observe from Tab. 1, Llama3.2 model itself is not strong enough to perform ideally on the reasoning tasks, so deviate from the reference model but not too much would give the model more freedom to have explorations to fetch higher rewards.
For the gradients as Fig. 4, we applied gradient normalization and clipping to stabilize training by limiting the total gradient norm to a fixed $\ell_{2}$-norm threshold for the whole model. But the visualized gradient is before clipping. We can see the KRPO has higher gradient than GRPO and gradually going down as the estimated variance from Kalman Filter process can highlight the benefits / drawbacks of actions. The optimization signal is stronger and more informative, making it easier to be optimized than its baseline model.
@@ 5.4.4 Case Study
We also provide a case study from OpenMath data as Tab. 2 to show the improved reasoning ability of KRPO model over GRPO baseline. From the case, we observe that GRPO can figure out the prior steps, e.g. “This implies $34x=$ 3, so $x=\frac{3}{4}$”, but it stops here and derive a wrong answer 343. On the contrary, the proposed KRPO can get the correct answer 1. For this question, the KRPO can get 12 corrects out of 12 trials, while GRPO can have 3 corrects out of 12. It is because KRPO can better award and punish the correct and incorrect answers via a more accurate advantage estimation.
@@ Appendix C The Analysis of Leading Statement for Language Model Reasoning
Interestingly, by enhancing with the leading statement (denoted with KRPO*), on the hard question set, the KRPO performance has increased greatly (from 20.881% of KRPO to 40.554% of KRPO*). However, for easy and normal question sets, there are improvements with leading statements but this advantage is not as obvious as on hard question set: on easy set, KRPO 71.764% V.S. KRPO* 72.478%; on the normal set, KRPO 50.589% V.S. 51.403%. It may be caused by the learning saturation of model on easy and normal question sets as the questions have fewer terms and digits that are not that difficult to be handled; but for hard question set with a lot more terms and digits, the model will easily get lost in the understanding of just arithmetic questions.

### v2

@@ 5.4.3 KL Divergence, Normalized Gradients and Reward Baseline Stabilization Analysis
(c) The Kalman filter gradually stabilizes the returns to a certain level.
@@
We also examine the KL divergence of action probabilities between the RL finetuned model and its reference model in Fig. 4(a). Interestingly, we notice that the KRPO has slightly higher KL divergence compared to GRPO baseline. It makes sense as we observe from Tab. 1, the Llama3.2 model itself is not strong enough to perform ideally on the reasoning tasks, so deviating from the reference model but not too much would give the model more freedom to explore and fetch higher rewards.
For the gradients as Fig. 4, we applied gradient normalization and clipping to stabilize training by limiting the total gradient norm to a fixed $\ell_{2}$-norm threshold for the whole model. But the visualized gradient is before clipping. We can see the KRPO has a higher gradient than GRPO and gradually goes down, as the estimated variance from the Kalman Filter process can highlight the benefits/drawbacks of actions. The optimization signal is stronger and more informative, making it easier to optimize than its baseline model.
@@ 5.4.4 Case Study
We also provide a case study from OpenMath data as Tab. 2 to show the improved reasoning ability of KRPO model over GRPO baseline. From the case, we observe that GRPO can figure out the prior steps, e.g. “This implies $34x=$ 3, so $x=\frac{3}{4}$”, but it stops here and derives the wrong answer 343. On the contrary, the proposed KRPO can get the correct answer 1. For this question, the KRPO can get 12 corrects out of 12 trials, while GRPO can have 3 corrects out of 12. This is because KRPO can better award and punish correct and incorrect answers via a more accurate advantage estimation.
@@ Appendix C The Analysis of Leading Statement for Language Model Reasoning
Interestingly, by enhancing with the leading statement (denoted with KRPO*), on the hard question set, the KRPO performance has increased largely (from 20.881% of KRPO to 40.554% of KRPO*). However, for easy and normal question sets, there are improvements with leading statements but this advantage is not as obvious as on hard question set: on easy set, KRPO 71.764% V.S. KRPO* 72.478%; on the normal set, KRPO 50.589% V.S. 51.403%. It may be caused by the learning saturation of model on easy and normal question sets as the questions have fewer terms and digits that are not that difficult to be handled; but for hard question set with a lot more terms and digits, the model will easily get lost in the understanding of just arithmetic questions, so some leading words can help greatly.